7) (10 pts) Find the principle values of the natural logarithm of 2 (1+1). Express in the form a + b
Answer #1

Here z= 9 = Hoj (1+i) (1+3)² = 14201 + j2 = 21 (1+339 = (23)= 43° = -4 ie, (1+3) Now log 640i) = log1+3.0 Since log (a+jb) =